# Sequential Causal Normal Form Games: Theory, Computation, and Strategic Signaling

### v1

@@ Introduction
Strategic interactions often unfold sequentially, with leaders committing to actions before followers respond. While classical Stackelberg games (Stackelberg1934) model such dynamics under perfect rationality, real-world agents exhibit biases, follow instincts, or engage in counterfactual reasoning that deviates from game-theoretic prescriptions (wason1974dual; sloman1996empirical).
Causal Normal Form Games (CNFGs) (Maiti2025) address this by incorporating the Pearl Causal Hierarchy (PCH) (Pearl2009): observational ($L_{1}$, instinctive), interventional ($L_{2}$, deliberate), and counterfactual ($L_{3}$, sophisticated reasoning). However, CNFGs are limited to simultaneous-move settings, while many strategic interactions are inherently sequential (letchford2010computing).
@@
CNFGs (Maiti2025) model multi-agent interactions through Structural Causal Models (SCMs) (Pearl2009), enabling agents to act across three layers of the PCH (Bareinboim2022):
@@
The key insight: different causal structures yield identical $L_{2}$ payoff matrices but different $L_{1}$ and $L_{3}$ outcomes, resolving paradoxes where ”irrational” behavior outperforms game-theoretic prescriptions (howard1971paradoxes; basu1994traveler).
@@
Bayesian Stackelberg games (Conitzer2006) model uncertainty over follower types using belief distributions but assume all agents operate at the interventional ($L_{2}$) level. Our framework allows heterogeneity in causal reasoning layers. Epistemic game theory (Perea2012) uses hierarchical belief models to capture strategic reasoning but does not explicitly model causal mechanisms or the PCH. Agents’ beliefs about others’ rationality differ from beliefs about causal structures. Psychological games (Geanakoplos1989) incorporate belief-dependent utilities but lack formal causal semantics. Our framework provides precise causal interpretations via SCMs. Signaling games (Spence1973; Cho1987) study information transmission but typically assume rational signaling ($L_{2}$). We show leaders can signal via causal layer choice, creating richer strategic possibilities. Causal influence diagrams (Hammond2023; Mishra2024) characterize interventions in simultaneous causal games. We extend this to sequential settings with timing constraints and information structures. Commitment in extensive-form games (letchford2010computing) studies commitment value in sequential games but under perfect rationality. We incorporate bounded rationality via PCH layers.
@@
  • Works for human agents (Simon1955; Simon1957; Kahneman1982; Rabin1993; Kaufmann1999; Toulis2016) (with cultural norms, emotions, bounded computation)
  • Works for classical economic agents (vonNeumann1944; Savage1954; Fishburn1989) (with well-defined utility functions)
  • Fails for AI agents (Fan2023; Richens2024; Jia2025; Zhang2024; Duan2024; Xu2023; Aher2024) whose ”instincts” (LLM priors) and ”reasoning” (inference procedures) don’t map cleanly to rational choice

### v2

@@ Introduction
Strategic interactions often unfold sequentially, with leaders committing to actions before followers respond. While classical Stackelberg games (von Stackelberg 1934) model such dynamics under perfect rationality, real-world agents exhibit biases, follow instincts, or engage in counterfactual reasoning that deviates from game-theoretic prescriptions (Wason and Evans 1974; Sloman 1996).
Causal Normal Form Games (CNFGs) (Maiti et al. 2025) address this by incorporating the Pearl Causal Hierarchy (PCH) (Pearl 2009): observational ($L_{1}$, instinctive), interventional ($L_{2}$, deliberate), and counterfactual ($L_{3}$, sophisticated reasoning). However, CNFGs are limited to simultaneous-move settings, while many strategic interactions are inherently sequential (Letchford and Conitzer 2010).
@@
CNFGs (Maiti et al. 2025) model multi-agent interactions through Structural Causal Models (SCMs) (Pearl 2009), enabling agents to act across three layers of the PCH (Bareinboim et al. 2022):
@@
The key insight: different causal structures yield identical $L_{2}$ payoff matrices but different $L_{1}$ and $L_{3}$ outcomes, resolving paradoxes where ”irrational” behavior outperforms game-theoretic prescriptions (Howard 1971; Basu 1994).
@@
Bayesian Stackelberg games (Conitzer and Sandholm 2006) model uncertainty over follower types using belief distributions but assume all agents operate at the interventional ($L_{2}$) level. Our framework allows heterogeneity in causal reasoning layers. Epistemic game theory (Perea 2012) uses hierarchical belief models to capture strategic reasoning but does not explicitly model causal mechanisms or the PCH. Agents’ beliefs about others’ rationality differ from beliefs about causal structures. Psychological games (Geanakoplos et al. 1989) incorporate belief-dependent utilities but lack formal causal semantics. Our framework provides precise causal interpretations via SCMs. Signaling games (Spence 1973; Cho and Kreps 1987) study information transmission but typically assume rational signaling ($L_{2}$). We show leaders can signal via causal layer choice, creating richer strategic possibilities. Causal influence diagrams (Hammond et al. 2023; Mishra et al. 2024) characterize interventions in simultaneous causal games. We extend this to sequential settings with timing constraints and information structures. Commitment in extensive-form games (Letchford and Conitzer 2010) studies commitment value in sequential games but under perfect rationality. We incorporate bounded rationality via PCH layers.
@@
  • Works for human agents (Simon 1955, 1957; Kahneman et al. 1982; Rabin 1993; Kaufmann 1999; Toulis and Parkes 2016) (with cultural norms, emotions, bounded computation)
  • Works for classical economic agents (von Neumann and Morgenstern 1944; Savage 1954; Fishburn 1989) (with well-defined utility functions)
  • Fails for AI agents (Fan et al. 2023; Richens and Everitt 2024; Jia et al. 2025; Zhang et al. 2024; Duan et al. 2024; Xu et al. 2023; Aher et al. 2024) whose ”instincts” (LLM priors) and ”reasoning” (inference procedures) don’t map cleanly to rational choice